OpenStudy (anonymous):

Write the solution in lowest terms. -1/3 + 3/4 =

OpenStudy (anonymous):

Options 5/12 2/12 1/6

OpenStudy (anonymous):

\[-1/3\times4/4+3/4\times3/4=-4/12+9/12=5/12\]